A monopolist earns $50 million annually and will maintain that level of profit indefinitely, provided no other firm enters the market. If another firm successfully enters the market, the incumbent's profits remain at $50 million the first period, but fall to $25 million annually thereafter. The opportunity cost of funds is 10 percent, and profits in each period are realized at the beginning of each period. What is the present value of the firm's current and future earnings if entry occurs?
Answer #1

Value of the indefinite earnings of $25 m = Cash flow/Rate

= 25m/10%

= $ 250 m

Present value of the earnings = C0+ CF1/(1+r)^1

= 50 + 250/1.1

= 50+ 227.27

= $ 277.27 million